Copyright\begin{matrix}c \\ ◯ \end{matrix}1999 ONG TCV Scoala Virtuala a Tanarului Matematician http://math.ournet.md 1

Модуль

1. |a| ≥ a.

2. |a| ≥ −a.

3. |a| = | − a|.

4. |a| ≥ 0 и |a| = 0 если и только если a = 0.

5. |a| = a если и только если a ≥ 0.

6. |a| = −a если и только если a ≤ 0.

7. |a| = \sqrt{a^{2}}.

8. |a + b| ≤ |a| + |b| и |a + b| = |a| + |b| если и только если ab ≥ 0.

9. |a · b| = |a| · |b|.

10. |\frac{a}{b}| = \frac{|a|}{|b|}.

11. |a| = b, b ≥ 0 ⇔ a = ±b ⇔ [ \begin{matrix}a = b, \\ a = −b. \end{matrix}

12. |a| < b, b > 0 ⇔ −b < a < b ⇔ [ \begin{matrix}a < b, \\ a > −b. \end{matrix}

13. |a| > b, b > 0 ⇔ [ \begin{matrix}a > b, \\ a < −b. \end{matrix}

14. |a| = |b| ⇔ (a − b)(a + b) = 0 ⇔ a = ±b.

15. |a| > |b| ⇔ (a − b)(a + b) > 0 ⇔ a ∈ (−∞; −|b|) ∪ (|b|; +∞) ⇔ −|a| < b < |a|.

16. |a|^{2} = a^{2}.